# Revisiting Entropy in Reinforcement Learning for Large Reasoning Models

### v1

@@ 4 Experimental Setup
We trained Qwen2.5-Math-7B (DBLP:journals/corr/abs-2409-12122) with RLVR using the veRL framework (DBLP:conf/eurosys/ShengZYWZZPL025) on the DAPO-Math-17K dataset (DBLP:journals/corr/abs-2503-14476). Training employed a rollout batch size of 256, generating 16 responses per prompt. The AdamW optimizer was applied with a cosine learning rate schedule and a peak rate of $1\times 10^{-6}$. During rollouts, decoding parameters were fixed as: top-$p=1.0$, temperature =1.0=1.0, and a maximum generation length of 4096 tokens.
@@
We evaluated both in-domain and out-of-domain performance. The in-domain benchmarks included AIME 2024/2025 (aime2024; aime2025), MATH500 (DBLP:conf/nips/HendrycksBKABTS21), AMC 2023 (amc2023), and Minerva Math (DBLP:conf/nips/LewkowyczADDMRS22), while out-of-domain evaluation covered LiveCodeBench (DBLP:conf/iclr/JainHGLYZWSSS25) for coding and IF-Eval (DBLP:journals/corr/abs-2311-07911) for instruction following. For each question, we sampled 64 responses and reported Avg@64 and Pass@64. Evaluation was performed under two decoding settings: (1) temperature 1.0, top-$p$ 1.0, and (2) temperature 0.6, top-$p$ 0.95, both with a maximum length of 4096 tokens. To mitigate potential data contamination in the Qwen2.5 series of LLMs (DBLP:journals/corr/abs-2507-10532), AIME 2025 served as the validation set, and the best checkpoint was evaluated on all other benchmarks.
5.1 Entropy and Response Diversity
@@
The diversity of responses generated by LLMs is strongly positively correlated with both the entropy of these responses and the entropy of the LLMs during training.
Figure 1: Evolution of entropy (solid lines) and N-gram diversity (markers) over training steps under different entropy regularization settings.
@@
To examine the relationship between response diversity and model entropy, we control LLM entropy using vanilla entropy regularization by varying the regularization coefficient. Experiments are conducted with coefficients $\{0.001,0.002\}$, without entropy regularization, and with a negative coefficient 0.001-0.001 that explicitly promotes entropy minimization. Response diversity is measured using the N-gram Diversity (DBLP:conf/naacl/LiGBGD16) and SelfBLEU (DBLP:conf/sigir/ZhuLZGZWY18) metrics. Formal definitions of these metrics are provided in Appendix A.
@@
Figure 1 shows the entropy dynamics during RLVR training alongside the N-gram Diversity of responses for AIME 2024 prompts. As shown in Figure 1, LLMs trained with GRPO exhibit entropy collapse during training, which is alleviated by vanilla entropy regularization with $\alpha=0.001$, while a larger coefficient ($\alpha=0.002$) causes entropy to increase continuously in later stages. Conversely, applying a negative coefficient ($\alpha=-0.001$) further exacerbates entropy collapse. Notably, N-gram Diversity follows a similar trajectory to entropy, with an average Pearson correlation of 0.8692 across the four training settings, indicating a strong positive correlation between response diversity and model entropy during training.
5.2 Entropy Dynamics on Prompts
@@
5.3 Prompt Entropy and Accuracy
@@
Figure 2: The ratio between the entropy of the LLMs on prompts during training and their entropy on the same prompts before training, plotted against training steps. Solid lines represent results on the AIME 2024, while dashed lines correspond to results on IF-Eval.
To examine the correlation between the entropy of LLMs on prompts and their corresponding accuracy, we measure the accuracy of each prompt as the proportion of correct responses among 64 generated responses. Figure 3 presents the scatter plot of entropy versus accuracy across AIME 2024, AIME 2025, and MATH500 for Qwen2.5-Math-7B. The figure shows that the entropy of LLMs on prompts is only weakly correlated with accuracy, as evidenced by the small average Pearson correlation coefficient (0.1542) across the three benchmarks. We further compute the Pearson correlation coefficient between entropy and accuracy at different training steps. As shown in Figure 4, although the coefficient fluctuates throughout training, it remains consistently small, confirming the weak correlation between entropy and accuracy.
5.4 Performance Gains Without Trading Off Entropy
To maintain the entropy of LLMs at a level comparable to that observed before training, we apply adaptive entropy regularization during RLVR training. Specifically, we randomly sample 1,000 prompts from the training dataset and compute the entropy of the LLMs on the corresponding responses prior to training. The computed entropy value is then set as the predefined threshold $\delta$, ensuring that the entropy of the LLMs remains above this threshold throughout training.
Figure 3: Scatter plot of accuracy versus entropy across AIME 2024, AIME 2025, and MATH500.
Figure 6 illustrates the evolution of entropy and accuracy on AIME 2024. As shown in Figure 6, under adaptive entropy regularization, the entropy of LLMs decreases sharply at the beginning of training, then increases, and subsequently fluctuates around the specified threshold. Meanwhile, accuracy on AIME 2024 exhibits a upward trend, even surpassing that of the model trained without adaptive entropy regularization. This observation indicates that performance improvement of LLMs is not merely achieved by trading off entropy. In contrast, LLMs trained without entropy regularization experience rapid entropy collapse, wherein entropy declines to a low value. Correspondingly, the Avg@64 metric rises sharply in the early stages of training but soon plateaus, remaining inferior to that of the model trained with entropy regularization. These results suggest that entropy collapse may lead to performance degradation.
Figure 4: Pearson correlation coefficients between the entropy of LLMs on prompts and their corresponding accuracy across different training steps. “Average” indicates the mean Pearson correlation coefficient computed over AIME 2024, AIME 2025, and MATH500.
5.5 Correlations Between Entropy and Model Performance
The correlations between LLM entropy and performance across benchmarks are task dependent. Specifically, the coding capabilities of LLMs exhibit a strong negative correlation with model entropy during training, whereas performance on tasks involving mathematical reasoning and instruction following shows relatively weak correlations with entropy.
The Pearson correlation coefficients between the entropy of LLMs and their performance across various benchmarks are presented in Table 1. As shown, when LLMs are trained with RLVR using only mathematical data, the Avg@64 and Pass@64 scores on LiveCodeBench exhibit strong negative correlations with model entropy during training, indicating that lower entropy is generally associated with stronger coding capabilities. This relationship is further illustrated in Figure 5, which shows a clear negative correlation between entropy and Avg@64 on LiveCodeBench. By contrast, the correlations between entropy and performance on other benchmarks are relatively weak.
Avg@64 Pass@64
AIME 2024 -0.23107 -0.10477
AIME 2025 -0.29374 0.10454
MATH500 0.11704 0.29805
AMC 2023 -0.18633 0.41696
Minerva -0.08230 0.38576
LiveCodeBench -0.79669 -0.71852
IF-Eval 0.67424 -0.61892
Table 1: Pearson correlation coefficients between the entropy of LLMs and their performance across different benchmarks. Coefficients with the largest absolute values are highlighted in bold.
@@ 5.5 Correlations Between Entropy and Model Performance
Figure 6: Evolution of the entropy of LLMs during training and their Avg@64 performance on AIME 2024 for models trained with RLVR, with and without adaptive entropy regularization. The predefined entropy threshold $\delta$ is set to 0.3657, corresponding to the entropy of the LLM on responses to 1,000 randomly sampled prompts from the training dataset prior to training. “Ada-Ent-Reg” denotes adaptive entropy regularization.
5.6 Entropy Collapse and Miscalibration
While RLVR enhances the performance of LLMs, it can lead to miscalibration, causing the models to become increasingly overconfident in their predictions. This miscalibration typically intensifies as entropy collapse becomes more severe. However, incorporating entropy regularization can mitigate this problem.
We further investigate the relationship between entropy and the calibration of LLMs. For a well-calibrated model, correct responses should receive higher probabilities, whereas incorrect ones should be assigned lower probabilities. To assess this, we compute the distributions of average per-token log probabilities for correct and incorrect responses generated by LLMs, as illustrated in Figure 7.
Figure 7: Distribution of log probabilities for correct and incorrect responses across different LLMs. “GRPO” denotes the LLM trained using GRPO, while “GRPO” followed by parentheses indicates LLMs trained with entropy regularization under varying coefficients. In each violin plot, the black line represents the mean.
As shown in Figure 7, prior to training, Qwen2.5-Math-7B generally assigns higher probabilities to correct responses than to incorrect ones. However, after GRPO training, the probabilities of both correct and incorrect responses increase, indicating that the model becomes more overconfident. Meanwhile, the probability gap between correct and incorrect responses narrows, suggesting reduced discriminability and poorer calibration, which aligns with the observations of DBLP:journals/corr/abs-2508-11800. Furthermore, when the entropy regularization coefficient is set to a negative value, thereby promoting entropy collapse, both overconfidence and miscalibration become more pronounced. In contrast, employing a positive entropy regularization coefficient mitigates these effects. Considering Figures 7 and 1 jointly, we observe a consistent trend in overconfidence and miscalibration: GRPO ($\alpha=-0.001$) > GRPO > GRPO ($\alpha=0.001$) > GRPO ($\alpha=0.002$). A similar ordering is observed for entropy collapse, suggesting a potential correlation between miscalibration and entropy collapse.
Previous studies have investigated the factors influencing the entropy of LLMs during training from both theoretical and empirical perspectives. From the theoretical perspective, for example, zhihu2025entropy and DBLP:journals/corr/abs-2505-22617 provided theoretical analyses identifying the factors that govern entropy dynamics during RLVR training. Specifically, they demonstrated that tokens exhibiting strong positive covariance between their probabilities and corresponding advantages primarily drive the substantial entropy reduction observed in RLVR. From the empirical perspective, DAPO (DBLP:journals/corr/abs-2503-14476) showed that raising the upper clipping threshold by increasing $\varepsilon_{\text{high}}$ can alleviate entropy collapse. However, the effect of $\varepsilon_{\text{low}}$, which controls the lower clipping threshold, on the entropy and performance of LLMs remains underexplored. Furthermore, empirical evidence from DBLP:journals/corr/abs-2505-22312 suggests that more off-policy updates exacerbate entropy collapse and degrade LLM performance on test sets. Nevertheless, how off-policy updates influence model performance on the training set remains unclear. To complement these findings, we conduct a series of experiments focusing on three aspects: (1) the effect of off-policy updates on entropy dynamics and model performance on both the training and test sets, (2) the influence of training data diversity on the entropy dynamics of LLMs, and (3) the impact of varying $\varepsilon_{\text{low}}$ on both model entropy and overall performance.
Figure 8: Evolution of the entropy and reward of LLMs during training.
6.1 Off-Policy Updates
More off-policy updates exacerbate entropy collapse, enabling the LLMs to achieve higher rewards on the training set, while their performance improvements on the test set are considerably less pronounced.
In GRPO, a batch of prompts is first sampled and used for rollout, advantage estimation, and the computation of log-probabilities with the LLM employed for rollout. This batch is then divided into several mini-batches, and the model parameters are updated once per mini-batch. Since the parameters change after the first mini-batch update, the data in the remaining mini-batches can be regarded as off-policy data with respect to the updated policy.
To examine the impact of off-policy data on the entropy and reward of LLMs during training, we varied the number of mini-batches per batch, which corresponds to the number of parameter updates per batch and is denoted as $N_{\text{update}}$, while keeping the rollout batch size and other hyperparameters fixed. Specifically, $N_{\text{update}}$ was set to 1, 2, and 4.
Figure 8 presents the evolution of entropy and reward under these configurations. As shown in the figure, increasing $N_{\text{update}}$ leads to a more pronounced entropy collapse, consistent with the findings of DBLP:journals/corr/abs-2505-22312, while yielding higher training rewards. Within the reward range of [0, 1], increasing $N_{\text{update}}$ from 1 to 4 raises the training reward by approximately 0.1 (equivalent to a 10% accuracy gain). However, despite this substantial improvement in training reward, Table 3 shows that the corresponding increase in the average Avg@64 score on in-domain test sets remains below 1%. Furthermore, as $N_{\text{update}}$ increases, the average Pass@64 score declines across both in-domain and out-of-domain benchmarks. These results suggest that more off-policy updates may cause the LLMs to overfit the training set.
6.2 Training Data Diversity
Lower data diversity leads to more pronounced entropy collapse during RLVR training. Moreover, training data size is not the only factor determining the performance of LLMs trained with RLVR, as LLMs trained on approximately 600 samples can perform comparably to those trained on about 17k samples.
To examine how training data diversity affects the entropy dynamics of LLMs, we train models on datasets with the same number of samples but different levels of diversity. Specifically, we construct subsets of the original training data using either K-means clustering or random sampling, ensuring that both methods select the same number of samples. Intuitively, subsets derived through K-means clustering are expected to exhibit lower diversity than those obtained via random sampling. Moreover, as the dataset size decreases, the overall data diversity is also expected to decline accordingly.
@@ 6.2 Training Data Diversity
Table 2: Performance of Qwen2.5-Math-7B trained with GRPO under varying data scales across multiple benchmarks. “Entropy” denotes the EMA-smoothed entropy at the final training step. Numbers in parentheses indicate the number of training samples, with subscripts “K-means” and “random” referring to datasets selected via K-means clustering and random sampling, respectively. “Average (ID)” and “Average (OOD)” denote the mean performance across in-domain and out-of-domain benchmarks. Results are reported as A / B, corresponding to Avg@64 and Pass@64, evaluated with temperature = 1.0, top-$p$ = 1.0, and a maximum generation length of 4096 tokens.
For subsets constructed using K-means clustering, we apply K-means with the number of clusters $K=1,000$. After clustering, the resulting clusters are sorted in descending order according to their sample counts. Each subset is then formed by selecting samples from the top $M$ clusters, where $M$ is set to 281, 112, 49, 21, and 9, corresponding to subsets containing 10,001, 5,031, 2,538, 1,246, and 616 samples, respectively. For subsets constructed via random sampling, we select the same numbers of samples (10,001, 5,031, 2,538, 1,246, and 616) as in the K-means based subsets to ensure a fair comparison. Finally, we train Qwen2.5-Math-7B with RLVR on both the K-means based and randomly sampled subsets, while keeping all training hyperparameters identical.

### v2

@@ Appendix B Experimental Setup
We trained Qwen2.5-Math-7B (Yang et al., 2024) with GRPO using the veRL framework (Sheng et al., 2025) on the DAPO-Math-17K dataset (Yu et al., 2025). Training employed a rollout batch size of 256, generating 16 responses per prompt. The AdamW optimizer was applied with a cosine learning rate schedule and a peak rate of $1\times 10^{-6}$. During rollouts, decoding parameters were fixed as: top-$p=1.0$, temperature =1.0=1.0, and a maximum generation length of 4096 tokens.
@@
We evaluated both in-domain and out-of-domain performance. The in-domain benchmarks included AIME 2024/2025 (MAA, 2024, 2025), MATH500 (Hendrycks et al., 2021), AMC 2023 (MAA, 2023), and Minerva Math (Lewkowycz et al., 2022), while out-of-domain evaluation covered LiveCodeBench (Jain et al., 2025) for coding and IF-Eval (Zhou et al., 2023) for instruction following. For each question, we sampled 64 responses and reported both Avg@64 and Pass@64. All evaluations used a decoding temperature of 1.0 and top-$p$ of 1.0. To mitigate potential data contamination in the Qwen2.5 series of LLMs (Wu et al., 2025), AIME 2025 served as the validation set, and the best checkpoint was evaluated on all other benchmarks.
Appendix C How Does the Entropy of LLMs Trained with RLVR Correlate with Their Performance?
C.1 Entropy and Response Diversity
@@
The diversity of responses generated by LLMs is strongly and positively correlated with their entropy during training.
@@
To examine the relationship between response diversity and model entropy, we control LLM entropy using vanilla entropy regularization by varying the regularization coefficient. Experiments are conducted with coefficients $\{0.001,0.002\}$, without entropy regularization, and with a negative coefficient 0.001-0.001 that explicitly promotes entropy minimization. Response diversity is evaluated using the N-gram Diversity (Li et al., 2016) and SelfBLEU (Zhu et al., 2018) metrics.
The N-gram Diversity metric (Li et al., 2016) measures the proportion of unique n-grams relative to the total number of n-grams in the generated responses. Let $U_{i}$ denote the number of unique n-grams and $C_{i}$ the total number of n-grams of order $i$. The metric is defined as follows:
$\text{N-gram Diversity}=\prod_{i=1}^{N}\frac{U_{i}}{C_{i}}.$ (8)
In our experiments, we set $N=5$ to assess response diversity.
SelfBLEU (Zhu et al., 2018) provides a complementary measure of diversity. For each response, the BLEU score is computed by treating it as the hypothesis and all other responses as references. The resulting BLEU scores, averaged over 1- to 4-gram, are then averaged across all responses to obtain the final SelfBLEU score.
@@
Figures 6 and 7 illustrate the entropy dynamics during RLVR training, together with the diversity of the generated responses to AIME 2024 prompts, as measured by N-gram Diversity and SelfBLEU, respectively. As shown in Figure 6, LLMs trained with GRPO exhibit entropy collapse during training, which is alleviated by vanilla entropy regularization with $\alpha=0.001$, while a larger coefficient ($\alpha=0.002$) causes entropy to increase continuously in later stages. Conversely, applying a negative coefficient ($\alpha=-0.001$) further exacerbates entropy collapse. Notably, N-gram Diversity follows a similar trajectory to entropy, with an average Spearman correlation of 0.8795 across the four training settings, indicating a strong positive correlation between response diversity and model entropy during training.
Figure 6: Evolution of entropy (solid lines) and N-gram diversity (markers) across training steps under different entropy regularization settings for Qwen2.5-Math-7B trained with GRPO.
C.2 Entropy Dynamics on Prompts
@@
Figure 7: Evolution of entropy (solid lines) and negative SelfBLEU (markers) across training steps under different entropy regularization settings for Qwen2.5-Math-7B trained with GRPO.
Figure 8: Ratio of the prompt entropy of Qwen2.5-Math-7B at various training steps relative to its initial entropy.
C.3 Prompt Entropy and Accuracy
@@
To examine the correlation between prompt entropy and response accuracy in LLMs, we quantify the accuracy of each prompt as the proportion of correct responses among 64 generated responses. Figure 9 shows scatter plots of prompt entropy versus accuracy on AIME 2024, AIME 2025, and MATH500 for Qwen2.5-Math-7B. The results show only a weak correlation, with an average Spearman’s rank coefficient of 0.0745 across the three benchmarks. We further compute the Spearman correlation at different training steps. As shown in Figure 10, although the coefficient fluctuates during training, it remains consistently small, further confirming the weak correlation between prompt entropy and response accuracy.
C.4 Correlations Between Entropy and Model Performance
The scatter plot depicting the relationship between LLM training entropy and model performance on LiveCodeBench, as measured by Avg@64, is presented in Figure 11.
Figure 9: Scatter plot of prompt entropy versus accuracy across AIME 2024, AIME 2025, and MATH500.
Appendix D What Factors Govern Entropy Dynamics, Both Theoretically and Empirically?
D.1 Clipping Threshold
A detailed description of the GRPO clipping variants explored in our study is presented below:
Figure 10: Spearman’s rank correlation coefficients between the entropy of LLMs on prompts and their corresponding accuracy across different training steps. “Average” indicates the mean Spearman’s rank correlation coefficient computed over AIME 2024, AIME 2025, and MATH500.
@@ D.1 Clipping Threshold
Table 3: Performance of Qwen2.5-Math-7B trained with GRPO under varying data scales. “Entropy” denotes the EMA-smoothed entropy at the final training step. Numbers in parentheses indicate the number of training samples, with subscripts “K-means” and “random” referring to datasets selected via K-means clustering and random sampling, respectively. “Average (ID)” and “Average (OOD)” denote the mean performance across in-domain and out-of-domain benchmarks. Results are reported as A / B, corresponding to Avg@64 and Pass@64.
  • Clip-Higher (Yu et al., 2025) raises the upper clipping bound in the GRPO objective to reduce the proportion of low-probability tokens being clipped. This relaxation allows these tokens to increase their likelihoods more freely, thereby enhancing exploration and mitigating entropy collapse in LLMs. Following Yu et al. (2025), we set $\varepsilon_{\text{high}}=0.28$.
  • Clip-Tighter, in contrast to Clip-Lower, decreases $\varepsilon_{\text{low}}$, thereby raising the lower clipping bound in the GRPO objective. Consequently, low-probability tokens with negative advantages are more likely to be clipped, preventing excessive suppression of their probabilities and mitigating entropy collapse in LLMs. Concretely, while Clip-Lower increases $\varepsilon_{\text{low}}$ from 0.2 to 0.28 (+0.08), Clip-Tighter symmetrically decreases it by 0.08, yielding a final value of 0.12.
D.2 Off-Policy Updates
The evolution of entropy and reward on the training set during LLM training under the Clip-Higher setting is presented in Figure 12.
D.3 Training Data Diversity
For K-means clustering, each prompt is represented by a mean-pooled token embedding from the final layer of Qwen2.5-Math-7B. We perform K-means clustering with $K=1,000$ and sort the resulting clusters in descending order according to their sample counts. Subsets are constructed by selecting samples from the top $M\in\{281,112,49,21,9\}$ clusters, yielding subsets of 10,001, 5,031, 2,538, 1,246, and 616 samples, respectively. For random sampling, we select the same number of samples to ensure a fair comparison. Qwen2.5-Math-7B is then trained with RLVR on all subsets using identical hyperparameters.